OpenStudy (anonymous):

There are 200 questions on a 3 hr examination.Among these questions are 50 mathematics problems.It is suggested that twice as much time be spent on each maths problem as for each other question.How many minutes should be spent on mathematics problems.

OpenStudy (anonymous):

i got 70 minutes

OpenStudy (anonymous):

200/180=1.1 ,twice -2.2 ,2.2*50=110 total=180 minutes,so 180-110=70

OpenStudy (anonymous):

is it right?

OpenStudy (jack1):

x = time spent per maths question y = time spent per other question so 50x + 150y = 180 minutes (3 hrs) and x = 2y (as you spend twice as much time on a maths q) so solve simultaneously

OpenStudy (jack1):

so y = 0.72 minutes (approx 43 seconds ish) and x = 1.44 minutes (approx 1 minute and 26 seconds ish) so 50x = total time spent on maths q's = 72 minutes

OpenStudy (jack1):

so use simultaneous equations to solve